OpenStudy (anonymous):

I=∫∫xydxdy. x^2+y^2≤7

OpenStudy (experimentx):

\[ \int_{-\sqrt 7 }^{\sqrt 7} \int_{-\sqrt{7 - x^2}}^{\sqrt{7 - x^2}} xy \; dy dx\]

OpenStudy (experimentx):

In polar coordinates \[ \int_0^\sqrt7 \int_0^{2 \pi} r^3\sin \theta \cos \theta d\theta dr\]

OpenStudy (dan815):

yes

OpenStudy (dan815):

let her work it out she needs some practice setting up integrals

OpenStudy (anonymous):

you scared me. i tried to do it on my on and i jumped straight to polar. why r^3? probably a stupid question but i need to understand it not just find out the answer

OpenStudy (dan815):

because rdrdtheta = polar and x = rcostheta y=rsintheta

OpenStudy (dan815):